Swing Arm Mechanics

End Stop position monitoring:

Side Date PP WP Ambient T Remarks
DX   18884 5872   after first calibration unit alignment
DX 2014-11-30 18884 5862   overwrote the park position -> not parked
DX 2015-02-xx -- 5872    
DX 2015-12-11 18870 5854 3.8C Park position switch was replaced with mechanical one in between
         
SX 2015-12-11   18662    
SX 2016-02-24   18662 2.21C  
DX 2016-02-24   5854 2.21C  
SX 2016-05-15   18661 8.00C  
DX 2016-05-15   5854 8.00C  
SX 2017-04-28   18663 1.8C The working pos switches have been changed to mechanical
DX 2017-04-28   5854 1.8C The working pos switches have been changed to mechanical
DX 2017-12-02   5854 3.5C  
SX 2017-12-02   18664 3.5C  

Issue Tracking

  • 2016-02-24 SX Working Position Switch: The working position switch was not changing to set after reaching working position.
    • 2016-03-08 Investigation on the switch showed that it was not initialized properly. This is a known issue with the electronic switches. When after a power cycle the 24V is not applied fully but to low for some time the switch does not initialize and remain in a fail state (fast blinking LED). Investigation: The switch was tested manually (climbing up) and monitored on the web interface. No reaction seen. Opening the switch the LED was found blinking. Resolution: The 24V line was detached and attached again. The switch was properly initialized and working again.
2016-10-16 DX CalUnit vignetting: The DX CalUnit showed strong vignetting in the yellow pupil of the WFS. Re-alignment of the DX Cal-Unit was considered as solution. The DX Calunit tilt was rotated toward park position of ~3 mm. This relaxed the positioning of the secondary but did not remove the vignetting. Further investigation lead to the conclusion that one fiber was loose. Investigation: Taking picture into the CalUnit on horizon. The picture show different illumination of the holograms. Resolution: Fixing loose fiber on the back of the CalUnit.

argos_dx_calunit_loose_fiber.jpg

Left side: In the yellow circle, the not fully illuminated hologrom due to the loose and tilted fiber. Right side: The same with the fiber fixed

Swing Arm Electronics

Known Issues:
  • Setup router.
  • Change the power lamp on front panel.
Malfunctions:
  • Green laser source had connection errors. Low temperature problem of the LED board?
  • White light source didn't connect and work. Low temperature problem?

Calibration Unit Mechanics

Open issues:
  • Test of fiber plate with bigger holes (0.2mm). Pre-aligned test plate is already at LBT.

Calibration Unit Optics

Tenon

Calibration Unit Verification

Verification Matrix

Calibration Unit Alignment

Side Date X Y Z RXSorted ascending RY Comment
DX 2015-02-21 -2.52 -3.77 4.35 -471 56  
DX 2015-05-01 -1.00 -6.50 -2.10 -310 215 night
DX 2015-04-30 -2.94 -8.18 -2.08 -150 12 day
DX 2015-05-01 -2.94 -8.18 2.08 -44 62 day
DX 2015-04-29 -2.59 -8.84 -1.96 46.9 109.6 day
DX 2015-04-28 -2.77 -8.90 -2.00 51.0 103.0 after crash
SX 2015-12-10 4.87 -0.03 7.80 215 -57 Focus to be adjusted later in the run (Done 20151213)
Side Date X Y Z RX RY Comment
Topic revision: r8 - 09 Aug 2018, AndrewColson
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback